A copper hot -water tank is in the form of a cylinder with a hemispherical cap and stands on a flat base . The internal diameter of the cylinder is 525 mm and the overall internal height is 1200 mm .

Calculate : (a) the volume of the tank in m ^ 3 (b) the capacity of the tank in litres


1
Expert's answer
2021-12-09T10:14:52-0500
"R_{cyl}=R_{hemisphere}=R=\dfrac{525}{2}mm=0.2625m"

"H_{cyl}=1.2m-R=1.2m-0.2625m=0.9375m"

"V_{cyl}=\pi R^2H_{cyl}"

"V_{hemisphere}=\dfrac{1}{2}(\dfrac{4\pi}{3}R^3)"

"V_{tank}=V_{cyl}+V_{hemisphere}"

"V_{tank}=\pi R^2H_{cyl}+\dfrac{2\pi}{3}R^3"

"=\dfrac{\pi R^2}{3}(3H_{cyl}+2R)"

(a)

"V_{tank}=\dfrac{\pi (0.2625m)^2}{3}(3(0.9375m)+2(0.2625m))"

"\approx0.240828848\ m^3"

The volume of the tank is "0.240828848" m3.


(b)

The capacity of the tank is "240.829" litres.